Determine an equation AND length for the median from vertex C for the triangle with vertices:

C(5, 2) | A(‐3, 3) | B(2, ‐5)


1
Expert's answer
2021-01-28T00:34:58-0500

Solution:

Median: A median is the line segment drawn from the vertex to the midpoint of the opposite side.


As per the image below we have M as the mid point of side AB.





Mid Point(M) = ( "\frac{x1 + x2}{2}" , "\frac{y1 + y2}{2}" )


M = ("\frac{-3 + 2}{2}" ,"\frac{3 + (-5)}{2}" )


M = ("\frac{-1}{2}" ,"-1" )


Now we need to find the equation of the line CM


We have the formula to find the equation of the line passing through two points.


"y-y1=m(x-x1)" ------------------------ (1)


Where m = "\frac{y2 - y1}{x2 - x1}"


Calculate m


m = "\frac{-1-2}{-1/2 - 5}"


m = "\frac{6}{11}"


Plug the value of m and ("x1" ,"y1" ) into equation (1)


"y- y1 = m(x-x1)"


"y-2 =\frac{6}{11}(x-5)"


"y-2 = \frac{6x}{11} - \frac{30}{11}"


"y= 2+\frac{6x}{11} - \frac{30}{11}"


"y= \frac{6x}{11}- \frac{8}{11}" --------------------------(2)


Distance:

Distance formula = "\sqrt{(x1-x2)^2 + (y1-y2)^2}"

Distance between Point C and Point M


D = "\sqrt{(5+\frac{1}{2})^2 +(2+1)^2}"


D = "\sqrt{(\frac{11}{2})^2 + 3^2}"

D = "\sqrt{\frac{121}{4} + 9}"


D = "\sqrt{\frac{157}{4}}"

D = 6.26